• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 12:27:18 on September 9,1999

    Entry number 26715

    keyword=DAQ mods, etc.

    A few DAQ modifications and other things during this shutdown:
    1. New version of halog. It should break less often
    on SunOS and HP-UX.
    2. Problem of counting H+ pulses gated by H+ (it
    was 7.5 Hz, should be 15 Hz) was fixed by adding 50
    nsec delay (to both H+ & -) to avoid marginal timing.
    3. Disks on adaqs2: Bad disk data4 replaced. New disk
    data7 added. Note: this may affect the filtering operation.
    4. Put some copies of recent manuals for CERN software
    (paw, etc) in the counting room file cabinet. Help yourselves.